Grade 12th passMagical Mathematics[Interesting Approach]

A police sees a thief at a distance of 600 Meyer.he chases him at a speed of 10km/hr and catches him in 12mins.find speed of thief in km/hr

Sanjay Kumar das
9 Years agoGrade 12th pass

1 Answer

Shubham gupta
9 Years ago
 Solved by:- Shubham gupta 
Let distance of thief =x
Distance of police= x+0.6 km,                                         Speed of police= 10km/hr.                               time  = 12 min or 1/5 hr.                                 distance = speed*time.                         Distance of police=10*1/5=2 km.               Distance of thief =x+0.6=2.          x= 1.4  ,       Speed of thief = distance /time= 1.4/0.2            = 7km/hr